Suppose a consumer has income of 200 birr per month and he wants to spend all of his income on two 

goods, X and Y, whose prices are birr 4 and birr 5 respectively. Based on this information answer the 

following questions:

A. Express the budget line of the consumer both algebraically and diagrammatically. 

B. Compute the equation of the budget line.

C. Determine the slope of the budget line and interpret the result


1
Expert's answer
2021-05-19T20:32:59-0400

a)

The budget line will be.

"M=P_xQ_x+P_yQ_y"


Diagrammaticaly.

b)


The equation of the budget line.

"M=P_xQ_x+P_yQ_y"

"Q_y=\frac{M}{P_y}-\frac{P_x}{P_y}\times\ Q_x"

"M=Income."

"P_y=Price\ of\ y"

"P_x=Price\ of\ x"

"Q_y=\frac{200}{5}-\frac{5}{5}x"

"Q_y=40-x"


c)

The slope of the budget line.

"Slope=-\frac{P_x}{P_y}"


"=\frac{0-40}{40-0}=\frac{-40}{40}"


"=-1"

The budget line has shifted to the left.